3. [lpt] Consider the wave pulses shown in snapshot graphs i-iv below. Each pulse has a time, absolute phase speed, and travel direction given below it. R) -3 -2 -1 1 23 45 6 7 X (n) 2 s, 1 m/s -3-2 -1 1 23 45 6 7 x (m) t1.5 , 2 m/s- iv) -1 -32 -1 1 23 45 6 7 x Cn) 1-4s, 1 m/s → -3 -2 -1 0 12 3 4 5 6 7 x (n)Select the corresponding history graph A-D given below (or none) for each snapshot graph. E.g.. if the fist snapshot has B as a comesponding history graph, the second, C, and the last two, none, enter BCNN. You only have 5 tries. A) B) 0 23 4 5 6 7 89 10 C) D) 3 -2 -11 2345 67 3 -2-1 1 234 5 6 7 Answer Submit All Answers
Answer #1

for

(i) Distance = Speed*Time

= 1*2 = 2 m

(ii) Distance = Speed*Time

= 2*1.5 = 3 m

(iii) Distance = Speed*Time

= 1*4 = 4 m

(iv) Distance = Speed*Time

= 1*3 = 3 m

Therefore the answer for (i) A, (ii) N, (iii) N, (iv) N

Therefore the overall answer which has to be fed in the space provided is ANNN